OpenStudy (fantageplayer):

Given f(x) = 4x + 13 and g(x) = 10x − 2, solve for (f + g)(x) and select the correct answer below. (f + g)(x) = 14x + 15 (f + g)(x) = 14x + 11 (f + g)(x) = 6x + 11 (f + g)(x) = 6x + 15

OpenStudy (fantageplayer):

@FaiqRaees

jimthompson5910 (jim_thompson5910):

`(f + g)(x)` is the same as `f(x)+g(x)`

OpenStudy (fantageplayer):

im confused

OpenStudy (fantageplayer):

i did another one like this but they gave me a value to plug in for x

jimthompson5910 (jim_thompson5910):

example: f(x) = 9x+5 g(x) = 7x-2 (f+g)(x) = f(x)+g(x) (f+g)(x) = (9x+5)+(7x-2) (f+g)(x) = 9x+5+7x-2 (f+g)(x) = 9x+7x+5-2 (f+g)(x) = 16x+3

OpenStudy (fantageplayer):

and this one dosent

OpenStudy (fantageplayer):

ohhhhhhhhhhhhhhhhhh

OpenStudy (fantageplayer):

(4x+13)+(10x+−2) =(4x+10x)+(13+−2) =14x+11 =14x+11?

jimthompson5910 (jim_thompson5910):

very good, you are correct

OpenStudy (fantageplayer):

can you help with one more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (fantageplayer):

Choose the polynomial written in standard form. x4y2 + 4x3y + 10x2 xy2 + 4x4y + 10x x4y2 + 4x3y5 + 10x2 x6y2 + 4x3y8 + 10x

jimthompson5910 (jim_thompson5910):

Let's focus on ` x4y2 + 4x3y + 10x2` first the first term of that expression is x^4y^2 the degree of this monomial is equal to the sum of the exponents, so 4+2 = 6 what is the degree of the next term 4x^3y ? hint: y = y^1

OpenStudy (fantageplayer):

7?

jimthompson5910 (jim_thompson5910):

4 is not an exponent in 4x^3y

OpenStudy (fantageplayer):

OHHHH

OpenStudy (fantageplayer):

3?

jimthompson5910 (jim_thompson5910):

4x^3y is the same as 4x^3y^1

jimthompson5910 (jim_thompson5910):

The exponents for 4x^3y^1 are 3 and 1 3+1 = 4 is the degree of 4x^3y^1

OpenStudy (fantageplayer):

oh

jimthompson5910 (jim_thompson5910):

what is the degree of the last term 10x^2 ?

OpenStudy (fantageplayer):

how does findng these degrees help determine if this expression is in standard form?

OpenStudy (fantageplayer):

and 3

jimthompson5910 (jim_thompson5910):

how are you getting a degree of 3 for 10x^2 ?

jimthompson5910 (jim_thompson5910):

IF there was a y on it like 10x^2y, then yes it would be 3 since 2+1 = 3 but 10x^2 has no other term with exponents on it. The only exponent is 2, so the degree is simply 2

jimthompson5910 (jim_thompson5910):

`how does findng these degrees help determine if this expression is in standard form?` that's a good question. The answer to this question is that a polynomial is in `standard form` only when the terms are arranged in order where the degrees of each monomial is decreasing example: 5x^3 + 6x^2 + 10x + 2 the degrees step down 3, 2, 1, 0 this is NOT in standard form 6x^2+3x^3+10x because the degrees are 2,3,1 in that order

jimthompson5910 (jim_thompson5910):

Let's look at ` x6y2 + 4x3y8 + 10x` first term is x^6y^2 and it has a degree of 6+2 = 8 second term is 4x^3y^8 and it has a degree of 3+8 = 11 The degrees are NOT decreasing since we went from 8 to 11 So ` x6y2 + 4x3y8 + 10x` is NOT in standard form